OpenStudy (anonymous):

can someone show me how to simplify 3sqrt128?

OpenStudy (anonymous):

sqrt(a*b) = sqrt(a) * sqrt(b). So if you find some square number factors of 128, you can pull them out. For example sqrt(25 * 4) = sqrt(25) * sqrt(4) = 5 * 2. Do something similar for 3*sqrt(128).

OpenStudy (anonymous):

\[3\sqrt{128}=3\sqrt{64}\sqrt{2}\] when mutliplying roots just multiply the numbers inside. the key to find a perfect square which is a multiple of the number inside \[3\left( 8 \right)\sqrt{2}=24\sqrt{2}\]

OpenStudy (precal):